# Relative Pose Estimation for Nonholonomic Robot Formation with UWB-IO Measurements

### v2

@@ III-C Relative Localization to Leader Robot
We point out that the odometry information $\textbf{z}^{\mathcal{O}_{i}}_{i}(t)$ of leader in (17) can be obtained by integrating $v_{0,h}$, $v_{0,z}$ and $w_{0}$ with the following Assumption 2. Besides, $\mathbf{z}^{\mathcal{O}_{i}}_{i}(t)$ can also be transmitted between robots via communication [1].
@@
Note that Assumption 2 indicates that each robot has some prior knowledge of the leader’s movements, which is reasonable and can also be found in cooperation localization works [19, 1]. The convergence of estimator (15) and (16) are presented in the following theorem.
@@ III-D Distributed Formation tracking Controller
[…]
@@
[…]
@@
[…]
where $k1>0$, $k2>0$, $k3>0$ and $k4>0$ are the control gain. Note that only the estimated formation tracking errors (III-D) are adopoted in controller (24). To achieve the formation control task, the following assumption should be satisfied,
@@
It is worth to point out that the persistently exciting angle velocity $w_{0}$ is a common condition to achieve asymptotically tracking control for nonholonomic robots and can be found in related results [38][1]. Furthermore, the requirement for $w_{0}$ to meet the PE condition is only for the asymptotically tracking control of nonholonomic robot systems, whereas the convergence of the proposed estimators (13) or (17) does not require this condition.
To better illustrate the calculation process of algorithm data collection, collaborative localization, and control instructions, we have added pseudocode for the proposed method in Appendix A-B. The stability of the proposed controller can be given in the following theorem.
@@ IV-A Performance of the PE-free relative localization method
Firstly, we evaluated the noisy robustness of the proposed method and the latest method [14]. Under the two-robot relative localization case, we add standard deviations ranging from 0.01m to 1m to each measurement of distance and odometer displacement. For each noise case, we set up 100 sets of simulation scenarios with different initial conditions. The parameters used in the simulation are $\Delta t=0.05$, $r_{0}=2$, $r_{0}=3$, $c_{0,w}=0.1$, $c_{1,w}=0.3$, $c_{0,v}=0.2$, $c_{1,v}=0.3$, $k_{1}=1$, $k_{2}=0.5$, $k_{3}=0.4$, $k_{4}=0.2$. For more information about parameter selection, please refer to Appendix B-A. Parameters adopted in the contrast simulation can be found in [14]. The simulation results are shown in Fig. 3. It can be seen that as the measurement noise level increases, the localization accuracy of both algorithms decreases. However, the performance of the proposed data-driven algorithm is still inferior to that of the PE condition based method in [14]. The fundamental reason is that the data-driven method can tolerate abnormal measurement values to a certain extent.
@@
Secondly, we evaluated the convergence speed and control performance of the proposed method and the latest method [14]. Under different swarm scales, we set up 100 sets of formation simulation scenarios with different initial conditions. The parameters used in the simulation are $\Delta t=0.05$, $r_{i}=2+2\mathrm{rand}(-0.5,0.5)$, $c_{i,w}=4\mathrm{rand}(-0.5,0.5)$, $\mathrm{rand}(0,1)$, $k_{1}=1$, $k_{2}=0.5$, $k_{3}=0.4$, $k_{4}=0.2$. Note that function $\mathrm{rand}(a,b)$ represents the random value from $a$ to $b$. In the comparative simulation, except for the localization and control algorithms, all other conditions were completely the same. The control performance is reflected by the smoothness of the motion trajectory. Define the smoothness indicator as:
$S_{i}=\int_{t_{0}}^{t_{\mathrm{end}}}\lVert v_{i}-v_{\mathrm{leader}}\rVert$ (25)
@@
where $t_{0}$ and $t_{\mathrm{end}}$ denote the beginning time and ending time of the simulation, $v_{i}$ is the velocity of robot $i$, $v_{\mathrm{leader}}$ is the velocity of the leader robot. The smoothness indicator(25) represents the degree to which the speed of robot $i$ deviates from the expected speed (i.e. leader speed). According to this definition, the convergence time and smoothness indicator(25) of the two algorithms is shown in Fig. 4 and Fig. 5. It can be seen that our algorithm generally has better convergence speed than existing results in swarms of different sizes. This is because the data-driven localization algorithm adopted both historical information and current information, which accelerates the convergence speed. Furthermore, one can see that the proposed method outperforms existing method [14] based on PE condition in terms of trajectory smoothness across various swarm sizes. This is because the PE condition based localization method [14] requires the relative displacement between robots satisfy the PE condition and add noise and other PE signals to the speed command. When the swarm size is 5, in a simulation experiment, the robot motion trajectories of two algorithms are shown in Appendix Appendix B-B. It can be clearly seen that our algorithm outperforms the PE based algorithm in terms of control performance (trajectory smoothness).
@@ IV-B Outlier measurement value detection
To reduce the impact of measurement outliers, an outlier detection method is proposed in Appendix A-B. The core idea of the detection method is that for any nominal distance measurement $d_{ij}[k]$, the absolute value of the difference between $d_{ij}[k]$ and any historical distance $d_{ij}[m]$ should be less than the sum of the robot’s displacement between the two measurements, i.e. $\lVert z_{i}^{\mathcal{O}_{i}}[k]-z_{i}^{\mathcal{O}_{i}}[m]\rVert+\lVert z_{j}^{\mathcal{O}_{j}}[k]-z_{j}^{\mathcal{O}_{j}}[m]\rVert$, which is actually a triangular inequality. 100 sets of relative localization experiments under different initial conditions are conducted to verify the effectiveness of the outlier detection trick. To simulate the actual situation, we added measurement noise with a standard deviation of 0.05 to the UWB and IO measurements. The UWB sensor has a certain probability of detecting outlier values each time, that is, measurement noise with a standard deviation of 3. The probabilities of outlier values occurring are 0.05, 0.1, 0.2, 0.4, 0.6, and 0.8, respectively. The parameters adopted are the same as those in the two-robot case in Section IV-A. The pose estimation error with and without outlier detection algorithms is shown in Fig. 6. The accuracy of outlier detection is shown in Fig. 7. From the simulation results, one can see that the outlier detection algorithms can effectively detect measurement outliers and improve the accuracy of localization algorithms.
Due to space limitations, we provide more simulation results in Appendix B-(C-E) of the extended version of this article, including validation of the effectiveness of the algorithm with a large scale of swarm, noise robustness, and sensor failures.
@@ V-A Experiment Setup
In Fig. 8, the diagram of the experimental system is illustrated. The platform for 3D experiments is our unmanned aerial vehicle built based on a PX4 flight controller. A kind of nonholonomic robot Turtlebot3 is adopted as the experimental platform for the 2D experiments. Each robot is equipped with an IO to measure its own displacement and a UWB sensor to measure the real-time distance to its neighbors. For the drone, a Visual inertial odometer (VIO) is adopted to improve the accuracy of IO. Note that VIO is a technically mature odometer that is also used in related distance and odometer based relative localization work [12, 10]. There are several threads running on a ground working station, each thread $i$ receives the real-time measurement information sent by robot $i$. Distributed communication between threads are imitated with robot operating system (ROS), the communication topology is also defined in Fig. 8. The proposed distributed relative localization and distributed formation controller for robot $i$ are also executed by thread $i$. It should be noted that despite the usage of the ground station in the experiments, the algorithm still runs in a distributed manner.
@@ V-B Grounded robot Formation Control Experiment
Multi-robot formation results are provided to verify the effectiveness of the localization and control method in 2D situation. In this experiment, $3$ follower robots and a leader robot are assigned to achieve the formation task with the prescribed relative position $\textbf{p}_{1}^{*}=[0.5,-0.5,0]^{T}$, $\textbf{p}_{2}^{*}=[-0.5,-0.5,0]^{T}$ and $\textbf{p}_{3}^{*}=[0,0.5,0]^{T}$ in leader odometer frame $\mathcal{O}_{i}$. Parameter used in the experiment are $\Delta t=0.1$, $r_{i}=0.3,i=0,...,3$, $c_{0,w}=0.5$, $c_{1,w}=0.2$, $c_{2,w}=0.2$, $c_{3,w}=0.5$, $k_{1}=1$, $k_{2}=1$, $k_{3}=0.5$, $k_{4}=0$. After the data collection stage, the leader robot maintain a constant velocity $v_{0}$ and angle velocity $w_{0}$ to achieve a circular formation.
@@
To verify the consistency of the algorithm’s results under different initial conditions, we conducted Tb3 swarm formation experiments under 10 different initial conditions, with each experiment lasting 150 seconds. We recorded the average estimation and control errors throughout the experiment, as well as the relative pose estimation and tracking control error results, as shown in the Fig. 12. From the results, it can be seen that the experimental results at different initial positions have good consistency, and the differences in the results are mainly due to the possible variations in the measurement noise of the sensors in each experiment. The snapshot of experiments starting from three different initial conditions in ten sets of experiments can be found in Appendix B-F.
@@ V-C Aerial robot Tracking Control Experiment
In this experiment, follower robot $1$ is assigned to track the leader robot with a prescribed relative position $\textbf{p}_{1}^{*}\triangleq[1.2,1.2,0]^{T}$ in leader odometer frame $\mathcal{O}_{i}$. Parameter used in the experiment are $\Delta t=0.1$, $r_{0}=0.3$, $r_{1}=0.5$, $c_{0,w}=-0.5$, $c_{1,w}=0.4$, $k_{1}=0.4$, $k_{2}=0.8$, $k_{3}=0.5$, $k_{4}=0.5$. The condition for the end of the data collection stage is that the ratio of the minimum to maximum eigenvalues of the data collection matrix $S_{ij}$ is greater than the preset value of 0.1, which can usually be estimated through simulation.
Dual aerial robots tracking experiment results are provided to verify the effectiveness of the relative localization method in 3D situation. In this experiment, NOKOV motion capture system is used to record the ground truth pose of the leader and follower robots. The snapshots of the experiment process is shown in Fig. 13, the trajectory of the leader robot and the follower robot are described with orange and yellow curves in Fig. 13. The estimated error $\tilde{\textbf{e}}_{1}$ defined in (22) and tracking error $\textbf{e}_{1}$ defined in (10) are shown in Fig. 14 A and Fig. 14 B, respectively. One can see that when the measurement data is collected sufficiently, the estimation error and tracking error converge fast. This experiment verifies the effectiveness of the proposed localization algorithm in 3D case.
@@ V-D Outdoor Formation Control Experiment
To demonstrate the effectiveness of our algorithm in other environments (such as the corridor you mentioned), we tested the relative localization and formation control algorithm in a building courtyard corridor environment. The parameter adopted is the same as the indoor experiment. A snapshot of the experimental process is shown in Fig. 15. From the snapshot, one can see that the robot swarm implements relative localization based formation control task based on UWB and IO measurements, verifying the effectiveness of our method.
@@ VI Conclusion
This article has addressed the relative localization based distributed formation control for multi nonholonomic robots. The motivation arises from applications such as object transport in indoor and underground environments. First, we proposed a concurrent-learning based relative pose estimator, enabling inter-robot relative localization within local frame through only UWB distance and IO measurements. Second, to determine the relative pose between each robot and the leader robot, a cooperative localization strategy is introduced. Third, we devised the distributed formation controller based on the relative pose estimators designed in this article. With the proposed methods, distributed formation control for nonholonomic robots is achieved with only UWB distance and IO measurements. Both theoretical analysis and indoor experiments involving four robots were conducted to validate the effectiveness and adaptability of the proposed methods.
-A Proof of Theorem 1
@@
One can see that the norm of the normalized rotation matrix $\lVert\hat{\textbf{R}}^{\mathcal{O}_{i}}_{\mathcal{O}_{j}}(t)\rVert$ in (-B) is bounded. Under Assumption 1 and Theorem 1, $\lVert\tilde{\textbf{p}}^{\mathcal{O}_{i}}_{ij,0}(t)\rVert$ converges to zero asymptotically. Under the assuming convergence of $\lVert\tilde{\textbf{q}}^{\mathcal{O}_{j}}_{j0,0}\rVert$ for $j$ in layer $l$, $\lVert\tilde{\textbf{q}}^{\mathcal{O}_{i}}_{i0,0}(t)\rVert$ converges asymptotically. Similarly, assuming that for robot $i$ in layer $l$, the estimation $\lVert\tilde{\textbf{Q}}^{\mathcal{O}_{i}}_{\mathcal{O}_{0}}\rVert$ converges to zero asymptotically. Then for robot $i$ in layer $l+1$, it has

### v4

@@ III-C Relative Localization to Leader Robot
Note that $\textbf{z}^{\mathcal{O}_{i}}_{i}(t)$ in (17) can be obtained by integrating $v_{0,h}$, $v_{0,z}$ and $w_{0}$ with the following Assumption 2.
@@
Assumption 2 indicates that each robot is available to the leader’s odometry information. In practical applications, that information can be achieved through hop count propagation [19, 2]. We provide the pseudocode of the hop count propagation algorithm and simulation results to prove that the communication delay is acceptable in Appendix B-B. Besides, that information can also be estimated by designing distributed signal observers, as detailed in [5]. The convergence of the estimator (15) and (16) are presented in the following theorem.
@@ III-D Distributed Formation tracking Controller
Define $\tilde{e}_{i,x}=\hat{e}_{i,x}-e_{i,x}$, $\tilde{e}_{i,y}=\hat{e}_{i,y}-e_{i,y}$, $\tilde{e}_{i,z}=\hat{e}_{i,z}-e_{i,z}$, $\tilde{e}_{i,c}=\hat{e}_{i,c}-(1-\cos\theta_{\Sigma_{i}}^{\Sigma_{0}})$ and $\tilde{e}_{i,s}=\hat{e}_{i,s}-\sin\theta_{\Sigma_{i}}^{\Sigma_{0}}$, define estimation error as,
@@
$\left\{\begin{aligned} v_{i,h}&=r_{i}w_{i},\quad v_{i,z}=c_{i,v}\sin(c_{i,v}t)\\ w_{i}&=c_{i,w},\end{aligned}\right$. (23)
@@
$\left\{\begin{aligned} v_{i,h}&=v_{0,h}-k_{1}\hat{e}_{i,x}+k_{2}w_{0}\hat{e}_{i,y}\\ v_{i,z}&=v_{0,z}-k_{4}\hat{e}_{i,z},\quad w_{i}=w_{0}-k_{3}\hat{e}_{i,s}.\end{aligned}\right$. (24)
where $k_{1}>0$, $k_{2}>0$, $k_{3}>0$ and $k_{4}>0$ are the control gain. Note that only the estimated formation tracking errors (III-D) are adopoted in controller (24). To achieve the formation control task, the following assumption should be satisfied,
@@
It is worth to point out that the persistently exciting angle velocity $w_{0}$ is a common condition to achieve asymptotically tracking control for nonholonomic robots and can be found in related results [37][2]. Furthermore, the requirement for $w_{0}$ to meet the PE condition is only for the asymptotically tracking control of nonholonomic robot systems, whereas the convergence of the proposed estimators (13) or (17) does not require this condition. To better illustrate the calculation process of algorithm data collection, cooperative localization, and control instructions, we have added pseudocode for the proposed method in Appendix B-B. The stability of the proposed controller can be given in the following theorem.
@@ IV-A Performance of the PE-free relative localization method
To verify the robustness of our proposed relative localization method in the presence of noise, we conducted comparative evaluation with the latest method detailed in [14] in the 3D scenarios. In scenarios involving two robots, we systematically introduced standard deviations ranging from 0.01m to 1m into the measurements of distance and odometer displacement. For each noise case, we established 100 different simulation scenarios by varying initial conditions. The parameters used in these simulations are $\Delta t=0.05$, $r_{0}=2$, $r_{0}=3$, $c_{0,w}=0.1$, $c_{1,w}=0.3$, $c_{0,v}=0.2$, $c_{1,v}=0.3$, $k_{1}=1$, $k_{2}=0.5$, $k_{3}=0.4$, $k_{4}=0.2$. For further details on parameter selection and parameter sensitivity analysis, please refer to Appendix C-A. The parameters employed in the comparative simulations can be referenced from [14]. The simulation results are shown in Fig. 3, which reveals that as the measurement noise level increases, the localization accuracy of both algorithms declines. However, the performance of the PE condition based method in [14] is still inferior to that of the proposed data-driven algorithm. That is because the data-driven method can tolerate abnormal measurement values to a certain extent.
@@
Next, we examine the convergence speed and control performance of both our proposed method and the latest method [14] in the 3D scenarios. Across various swarm scales, we set up 100 formation simulation scenarios with different initial conditions. The parameters for these simulations are $\Delta t=0.05$, $r_{i}=2+2\mathrm{rand}(-0.5,0.5)$, $c_{i,w}=4\mathrm{rand}(-0.5,0.5)$, $\mathrm{rand}(0,1)$, $k_{1}=1$, $k_{2}=0.5$, $k_{3}=0.4$, $k_{4}=0.2$. The function $\mathrm{rand}(a,b)$ represents s random value between $a$ and $b$. In the comparative simulation, we maintained identical conditions except for the localization and control algorithms under analysis. The control performance is evaluated based on the smoothness of the motion trajectory, defined by the root-mean-square (RMS) velocity tracking error:
$S_{i}=\sqrt{\frac{1}{t_{\mathrm{end}-t_{0}}}\int_{t_{0}}^{t_{\mathrm{end}}}\lVert\bm{v}_{i}-\bm{v}_{\mathrm{leader}}\rVert^{2}\mathrm{d}t}$ (25)
@@
where $t_{0}$ and $t_{\mathrm{end}}$ denote the start and end time of the simulation, $\bm{v}_{i}$ and $\bm{v}_{\mathrm{leader}}$ are the velocity of robot $i$ and the leader robot. For first-order nonholonomic system, the smoothness indicator (25) quantifies how much the speed of robot $i$ deviates from the expected speed (i.e. leader speed). The convergence time and smoothness indicator (25) of the two algorithms are depicted in Fig. 4 and Fig. 5. Our results demonstrate that our algorithm consistently exhibits superior convergence speed across different swarm sizes, attributable to its incorporation of both historical and current information, thereby enhancing convergence efficiency. We also analyze the communication bandwidth required for algorithm deployment and discuss the time required for the algorithm to run on a laptop and an onboard computer (NVIDIA Jetson Orin). Please refer to the Appendix B-C for details. Our method outperforms the existing PE condition-based method from [14] in terms of trajectory smoothness across various swarm configurations, as the latter necessitates that robots execute PE localization enhancement motions to achieve relative localization, which adds noise and other PE signals to the speed command. When the swarm size is 5, in a simulation scenarios, the robot motion trajectories of two algorithms are shown in Appendix C-B. It can be clearly seen that our algorithm outperforms the PE based algorithm in terms of control performance.
@@ IV-B Outlier measurement value detection
To mitigate the impact of measurement outliers, an outlier detection method is proposed in Appendix B-B. The core idea of the detection method is based on the following principle: for any nominal distance measurement $d_{ij}[k]$, the absolute value of the difference between $d_{ij}[k]$ and any historical distance $d_{ij}[m]$ should be less than the sum of the robot’s displacements between the two measurements, i.e. $\lVert z_{i}^{\mathcal{O}_{i}}[k]-z_{i}^{\mathcal{O}_{i}}[m]\rVert+\lVert z_{j}^{\mathcal{O}_{j}}[k]-z_{j}^{\mathcal{O}_{j}}[m]\rVert$, which is actually a triangular inequality. To validate the outlier detection technique, we conduct 100 sets of relative localization experiments with varying initial conditions. To simulate the realistic situation, we introduce measurement noise with a standard deviation of 0.05 to the UWB and IO measurements. Additionally, the UWB sensor encounters outlier values with a measurement noise having a standard deviation of 3. The probabilities of detecting outlier values are set at 0.05, 0.1, 0.2, 0.4, 0.6, and 0.8, respectively. The parameters adopted are the same as those in the two-robot case presented in Section IV-A. Fig 6 and7 illustrate the pose estimation error with and without outlier detection algorithms, as well as the accuracy of outlier detection. The simulation results demonstrated that the outlier detection algorithms can effectively detect measurement outliers, thus enhancing the accuracy of localization algorithms. More simulation results are provided in Appendix C-(C-F), including validation of the effectiveness of the algorithm with dynamic topology robustness, a large scale of swarm, noise robustness, and sensor failures.
@@ V Experiment results
Both 2-D and 3-D localization and control experiments are provided to verify the effectiveness of the proposed method.
@@ V-A Experiment Setup
In Fig. 8, the diagram of the experimental system is illustrated. For the 3D experiments, we utilized an unmanned aerial vehicle based on a PX4 flight controller. For the 2D experiments, we employed a nonholonomic robot Turtlebot3 as the experimental platform. Each robot is equipped with an IO to measure its displacement and a UWB sensor to measure the real-time distance to its neighbors. For the aerial robot, a Visual inertial odometer (VIO) is adopted to improve the accuracy of IO. The VIO is a technically mature odometer also used in related distance and odometer based relative localization work [38, 11]. Several threads operate on a ground working station, with each thread $i$ receiving the real-time measurement information sent by robot $i$. Distributed communication between threads are imitated with robot operating system (ROS), and the communication topology is also defined in Fig. 8. The proposed distributed relative localization and distributed formation controller for robot $i$ are also executed by thread $i$. It is important to note that despite the ground station’s involvement in the experiments, the algorithm still operates in a distributed manner.
@@ V-B Grounded robot Formation Control Experiment
Multi-robot formation results are provided to verify the effectiveness of the localization and control method in 2D situation. In this experiment, $3$ follower robots and a leader robot are assigned to achieve the formation task with the prescribed relative position $\textbf{p}_{1}^{*}=[0.5,-0.5,0]^{T}$, $\textbf{p}_{2}^{*}=[-0.5,-0.5,0]^{T}$ and $\textbf{p}_{3}^{*}=[0,0.5,0]^{T}$ in leader odometer frame $\mathcal{O}_{0}$. Parameter used in the experiment are $\Delta t=0.1$, $r_{i}=0.3,i=0,...,3$, $c_{0,w}=0.5$, $c_{1,w}=0.2$, $c_{2,w}=0.2$, $c_{3,w}=0.5$, $k_{1}=1$, $k_{2}=1$, $k_{3}=0.5$, $k_{4}=0$. After the data collection stage, the leader robot maintain a constant velocity $v_{0}$ and angle velocity $w_{0}$ to achieve a circular formation.
@@
To verify the consistency of the algorithm’s results under various initial conditions, we conducted Tb3 swarm formation experiments under 10 different initial conditions, with each experiment lasting 150 seconds. We recorded the average estimation and control errors throughout the experiment, as well as the relative pose estimation and tracking control error results, as shown in Fig. 12. The results indicate that the experimental results at different initial positions exhibit strong consistency. A snapshot of experiments, initiated from three different initial conditions in ten sets of experiments, can be found in Appendix C-G.
@@ V-C Aerial robot Tracking Control Experiment
In this experiment, the follower robot $1$ is assigned with tracking the leader robot, maintaining a prescribed relative position $\textbf{p}_{1}^{*}\triangleq[1.2,1.2,0]^{T}$ in leader odometer frame $\mathcal{O}_{i}$. The experimental parameters are $\Delta t=0.1$, $r_{0}=0.3$, $r_{1}=0.5$, $c_{0,w}=-0.5$, $c_{1,w}=0.4$, $k_{1}=0.4$, $k_{2}=0.8$, $k_{3}=0.5$, $k_{4}=0.5$. The data collection stage ends when the ratio of the minimum to maximum eigenvalues of the data collection matrix $S_{ij}$ exceeds a preset threshold of 0.1, which can typically be estimated via simulation.
Results from the dual UAV tracking experiment demonstrate the effectiveness of the relative localization method in a 3D context. In this experiment, a NOKOV motion capture system is used to record the ground truth pose of the leader and follower robots. Snapshots of the experiment process are shown in Fig. 13 where the trajectories of the leader robot and the follower robot are presented with orange and yellow curves. The estimated error $\tilde{\textbf{e}}_{1}$ defined in (22) and tracking error $\textbf{e}_{1}$ defined in (10) are shown in Fig. 14 A and Fig. 14 B, respectively. The results indicate that with sufficient data collection, both the estimated error and tracking error converge rapidly. This outcome affirms the effectiveness of the proposed localization algorithm in the 3D tracking scenario.
@@ V-D Outdoor Formation Control Experiment
To further demonstrate the effectiveness of our algorithm in other environments, we tested the relative localization and formation control algorithm in a building courtyard corridor setting. The parameters adopted are consistent with those used in the indoor experiment. A snapshot of the experimental process is shown in Fig. 15. As depicted, the robot swarm effectively execute the formation control task based on UWB and IO measurements, verifying the effectiveness of our method. For more complex outdoor environments (such as foggy environments, rainy environments, etc.), UWB-IO fusion based localization systems can be studied in the future.
@@ VI Conclusion
This article has addressed the relative localization based distributed formation control for multi nonholonomic robots. The motivation arises from applications such as object transport in indoor and underground environments. First, we proposed a concurrent-learning based relative pose estimator, enabling inter-robot relative localization within local frame through only UWB distance and IO measurements. Second, to determine the relative pose between each robot and the leader robot, a cooperative localization strategy is introduced. Third, we devised the distributed formation controller based on the relative pose estimators designed in this article. With the proposed methods, distributed formation control for nonholonomic robots is achieved with only UWB distance and IO measurements. Both theoretical analysis and indoor experiments involving four robots are conducted to validate the effectiveness and adaptability of the proposed methods.
Appendix A Theoretical analysis
A-A Proof of Theorem 1
@@ A-B Proof of Theorem 2
One can see that the norm of the normalized rotation matrix $\lVert\hat{\textbf{R}}^{\mathcal{O}_{i}}_{\mathcal{O}_{j}}(t)\rVert$ in (A-B) is bounded. Under Assumption 1 and Theorem 1, $\lVert\tilde{\textbf{p}}^{\mathcal{O}_{i}}_{ij,0}(t)\rVert$ converges to zero asymptotically. Under the assuming convergence of $\lVert\tilde{\textbf{q}}^{\mathcal{O}_{j}}_{j0,0}\rVert$ for $j$ in layer $l$, $\lVert\tilde{\textbf{q}}^{\mathcal{O}_{i}}_{i0,0}(t)\rVert$ converges asymptotically. Similarly, assuming that for robot $i$ in layer $l$, the estimation $\lVert\tilde{\textbf{Q}}^{\mathcal{O}_{i}}_{\mathcal{O}_{0}}\rVert$ converges to zero asymptotically. Then for robot $i$ in layer $l+1$, it has